# Determination of the initial condition for the Balitsky-Kovchegov equation with transformers

### v1

@@ Abstract
In the high-energy limit of QCD, scattering off nucleons and nuclei can be described in terms of Wilson-line correlators whose energy dependence is perturbative. The energy dependence of the two-point correlator, called the dipole amplitude, is governed by the Balitsky–Kovchegov (BK) equation. The initial condition for the BK equation can be fitted to the experimental data, which requires evolving the dipole amplitude for a large set of different parameter values. In this work, we train a transformer model to learn the energy dependence of the dipole amplitude, skipping the time-consuming numerical evaluation of the BK equation. The transformer predicts the learned dipole amplitude and the leading order inclusive deep inelastic scattering cross section very accurately, allowing for efficient fitting of the initial condition to the experimental data. Using this setup, we fit the initial condition of the BK equation to the inclusive deep inelastic scattering data from HERA and consider two different starting points $x_{0}$ for the evolution. We find better agreement with the experimental data for a smaller $x_{0}$. This work paves the way for future studies involving global fits of the dipole amplitude at leading order and beyond.
@@ 1 Introduction
Mapping the partonic structure of protons and nuclei in the high-energy regime is a foremost challenge that provides the primary scientific motivation for the future Electron-Ion Colliders Accardi:2012qut ; AbdulKhalek:2021gbh ; Anderle:2021wcy . At a very small Bjorken-$x$ variable, the density of gluons within a hadron grows rapidly, leading to the gluon saturation phenomenon where non-linear QCD dynamics become dominant. The Color Glass Condensate (CGC) effective field theory Iancu:2003xm ; Gelis:2010nm provides a powerful framework for describing this dense gluonic matter. Central to the CGC is the dipole–target scattering amplitude, which encodes the essential information about the nonperturbative scattering off the gluonic target. The energy evolution of the dipole amplitude is governed by perturbative evolution equations, most notably the Balitsky–Kovchegov (BK) equation Balitsky:1995ub ; Kovchegov:1999yj . However, the starting point for this evolution—the initial condition at a moderately small $x_{0}$—is fundamentally nonperturbative and must be determined from experimental data.
@@
A convenient initial condition for the high-energy evolution is given by the McLerran–Venugopalan (MV) model McLerran:1993ka ; McLerran:1993ni ; McLerran:1994vd , which allows for a description of the dipole amplitude in terms of a few free parameter that can be extracted from the experimental data. The MV model, and its generalizations, have then be used in many successful extractions at both leading order Albacete:2009fh ; Lappi:2013zma ; Albacete:2010sy ; Ducloue:2019jmy ; Beuf:2020dxl ; Dumitru:2023sjd ; Casuga:2023dcf ; Hanninen:2025iuv and next-to-leading order Hanninen:2022gje ; Casuga:2025etc using the inclusive deep inelastic scattering (DIS) data from HERA H1:2009pze ; H1:2015ubc . The extracted parameters contain the fundamental information about the target structure, including the saturation scale $Q_{s}$ that describes the onset of the gluon saturation phenomenon.
@@
However, to determine the initial condition from the experimental data, one has to run the BK evolution for a wide range of parameter values. Due to the nonlinear nature of the BK evolution, its evaluation is much more demanding than the linear BFKL Lipatov:1976zz ; Fadin:1975cb ; Kuraev:1976ge ; Kuraev:1977fs ; Balitsky:1978ic and DGLAP Dokshitzer:1977sg ; Gribov:1972ri ; Gribov:1972rt ; Altarelli:1977zs evolution equations, making the BK evolution the bottleneck for numerical studies even at leading order. At higher orders in perturbation theory, the cross section itself becomes also numerically demanding due to the large amount of integrals over transverse coordinates that are typical for the dipole picture used in small-$x$ calculations. For these reasons, it becomes important to study efficient methods for evaluating these quantities with different models for the dipole amplitude.
To overcome this limitation, we turn to the rapidly advancing field of machine learning and the neural network approach. We employ a large language model transformer architecture, a class of neural networks renowned for its ability to capture long-range correlations and complex patterns in sequential data. By treating both the model parameters and the kinematic variables as a sequence, the transformer’s self-attention mechanism can learn the intricate, non-local relationships within the experimental dataset (such as those from HERA) without preconceived notions of the underlying physics. This paper presents the first determination of the small-$x$ dipole amplitude using this novel transformer-based method. Moreover, we consider different starting points for the BK evolution, enabling us to study the bias in the initial condition and to assess the validity of the BK evolution for moderate values of the Bjorken-$x$ variable. It should be emphasized that the present analysis is restricted to leading-order accuracy, and our primary aim is to demonstrate the methodology rather than to obtain the most precise phenomenological fit. Gaussian-process emulators have been successfully applied in small-$x$ studies Casuga:2023dcf ; Casuga:2025etc , providing flexible surrogate models with quantified uncertainties. In this work, we employ a transformer-based approach, which can naturally accommodate larger training sets, capture correlations across the parameter space, and demonstrates stable performance even when modest extrapolation beyond the sampled region is required. This makes it well suited for global analyses requiring high precision over a wide kinematic range.
@@ 2.1 Inclusive DIS in the dipole picture
Inclusive DIS is described by the structure functions $F_{2}(x,Q^{2})$ and $F_{L}(x,Q^{2})$, which are related to the experimental reduced cross section $\sigma_{r}$ by
@@
In the small-$x$ regime, these cross sections can be calculated within the dipole picture Kovchegov:2012mbw , where the interaction factorizes into the splitting of the virtual photon into a quark–antiquark pair ($q\bar{q}$), followed by the scattering of this color dipole off the proton target. The total cross section is given by an integral over the dipole’s transverse size $\mathbf{r}$, its impact parameter $\mathbf{b}$, and the quark’s longitudinal momentum fraction $z$, which reads
@@
and the argument $\overline{Q}$ of the modified Bessel functions $K_{0}$ and $K_{1}$ is $\overline{Q}^{2}\equiv z(1-z)Q^{2}$. Here, we have ignored quark mass corrections.
@@
and introducing a single normalization parameter, $\sigma_{0}/2$, which has the interpretation as the proton transverse area  Lappi:2013zma ; Albacete:2010sy ; Albacete:2009fh . This parameter effectively sets the total strength of the interaction by replacing the impact parameter integral in the cross-section calculation. It is a free parameter in our model, constrained by the fit to HERA data. The integrated dipole amplitude $N(r,x)$ now only depends on the dipole size $|r=\absolutevalue{{\mathbf{r}}}$ and the Bjorken $x$ variable.
@@ 2.2 Balitsky–Kovchegov evolution
The energy dependence of the dipole amplitude, $N(r,x)$, is governed by the BK evolution equation. This equation resums quantum corrections proportional to large logarithms of energy, $\ln(1/x)$, and incorporates the non-linear effects that lead to gluon saturation. For a dipole with quark and antiquark at transverse positions ${\mathbf{r}}_{0}$ and ${\mathbf{r}}_{1}$, the equation reads
@@
The interaction kernel, $K_{\text{BK}}$, describes the splitting of the parent dipole. For phenomenological accuracy, it is crucial to include running coupling corrections. Following the Balitsky prescription Balitsky:2006wa , the kernel is given by
@@
Here the position-space strong coupling constant, $\alpha_{s}(r^{2})$, is taken at leading order with $n_{f}=3$ active quark flavors as Lappi:2013zma
@@
Here, $\Lambda_{\text{QCD}}=$0.241\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$\$ and $C^{2}$ is a parameter that sets the scale of the argument of the logarithm, which must be determined from fits to data Lappi:2013zma ; Casuga:2023dcf . For large values of the dipole size $r$, we freeze $\alpha_{s}$ to the value $\alpha_{s,\text{max}}=0.7$.
@@ 3.1 Sampling and solving the BK equation
The BK equation (8) describes the evolution of the dipole amplitude $N(r,x)$ with respect to the collision energy. The evolution is uniquely specified by an initial condition at $x_{0}$, and to assess the dependence on the initial rapidity scale we consider two different choices $x_{0}=0.01$ and $x_{0}=0.05$. For the initial condition, we use a generalization of the MV model considered in Ref. Lappi:2013zma
@@
which is a function of the initial saturation scale $Q_{s0}$, the anomalous dimension $\gamma$, and an infrared regulator $e_{c}$. To efficiently map this parameter space, we use Latin hypercube sampling (LHS) to generate 10 00010\,000 parameter sets for $(Q_{s0},\gamma,e_{c},C^{2})$. For each set, the running-coupling BK equation with the Balitsky kernel is solved using a numerical Julia implementation developed for this work. The solution is tabulated on a two-dimensional grid spanning transverse sizes $10^{-6}\leq r/\mathrm{GeV}^{-1}\leq 10^{2}$ and rapidities $0\leq\ln(x_{0}/x)\leq 16$, which covers the relevant HERA kinematics H1:2015ubc . This procedure yields a dataset of approximately $4.7\times 10^{7}$ values of $N(r,x)$, which serves as the training data for the neural network model.
@@ 3.2 Neural network emulation of the BK amplitude
Figure 1: Left: Two dimensional histogram comparing the exact and transformer emulated values of $\ln N(r,x)$ for the BK dipole amplitude. The color scale indicates the density of samples, and the dashed line represents exact agreement. The mean and median relative errors on the validation set are $0.09\%$ and $0.05\%$, respectively. Right: Distribution of the signed relative errors, zoomed to $\pm 1\%$. This highlights that the vast majority of predictions fall within a few per mille of the true values, demonstrating the emulator’s high accuracy.
Directly solving the BK equation for every trial set of parameters in a fit would be computationally prohibitive. To circumvent this we train a neural network to emulate the functional dependence of $\ln N(r,x)$ on the variables $(r,x)$ and the three parameters of the initial condition. Owing to its ability to capture correlations in an ordered sequence of input variables, we adopt a transformer architecture. The network takes as input the 6-dimensional vector $(\log_{10}Q_{s0},\,e_{c},\,\gamma,\,\log_{10}C,\,\log_{10}(r/\mathrm{GeV}^{-1}),\,x)$. Each scalar is mapped by a shared linear projection to form a sequence of six tokens of dimension $d_{\rm model}=128$. A sinusoidal positional encoding is added to the sequence, which is then processed by a transformer encoder with four layers, eight attention heads, feed-forward width 512, dropout 0.1, and pre–normalization. The token representations are mean-pooled and passed to a two-layer multilayer perceptron (MLP) head GARDNER19982627 that outputs a raw logit function pmlr-v162-wei22d . Applying a sigmoid to this logit yields the dipole amplitude $N(r,x)$, automatically enforcing the unitarity bound $0\leq N\leq 1$. For diagnostics we also use $\log N$ after the sigmoid.
@@
Inputs are standardized using the StandardScaler implementation from the Scikit-learn library pedregosa2011scikit . The model is trained using the AdamW optimizer Loshchilov2017DecoupledWD with learning rate $3\times 10^{-4}$, weight decay $10^{-4}$, cosine annealing schedule to $3\times 10^{-6}$, batch size $4096$, and early stopping (patience $20$). Automatic mixed-precision training Micikevicius2017MixedPT and gradient clipping (max-norm $1.0$) are applied. $90\%$ of the generated data is used for training and the remainder for validation.
Figure 2: The emulated dipole amplitude $N(r,x)$ (lines) is compared against the exact BK solution (points) for a representative out-of-training-sample parameter set at three different rapidities, $x=10^{-3}$, $10^{-5}$, and $10^{-7}$. The excellent agreement in both log-linear (left) and log-log (right) scales across the entire range of $r$ demonstrates the model’s robust extrapolation capability. The chosen parameters are $Q_{s0}^{2}\simeq 0.07$ GeV2, $\gamma=1.01$, $e_{c}=24.68$, $C^{2}\simeq 4.65$, and $x_{0}=0.01$, corresponding roughly to physical values found in previous fits Lappi:2013zma ; Casuga:2023dcf .
The trained emulator exhibits excellent performance, achieving mean and median relative errors of $0.09\%$ and $0.05\%$, respectively, on the validation set. This high accuracy is visually demonstrated in Fig. 1. The left panel shows a strong one-to-one correlation between the emulated and exact values of $\ln N(r,x)$ across many orders of magnitude. The distribution of relative errors, shown in the right panel, is sharply peaked at zero, confirming that the vast majority of predictions are accurate to within a few per mille.

### v2

@@ Abstract
In the high-energy limit of QCD, scattering off nucleons and nuclei can be described in terms of Wilson-line correlators whose energy dependence is perturbative. The energy dependence of the two-point correlator, called the dipole amplitude, is governed by the Balitsky–Kovchegov (BK) equation. The initial condition for the BK equation can be fitted to the experimental data, which requires evolving the dipole amplitude for a large set of different parameter values. In this work, we train a transformer model to learn the energy dependence of the dipole amplitude, thereby replacing repeated time-consuming numerical BK evolutions during the fit by fast emulator predictions. The transformer predicts the learned dipole amplitude and the leading order inclusive deep inelastic scattering cross section very accurately, allowing for efficient fitting of the initial condition to the experimental data. Using this setup, we fit the initial condition of the BK equation to the inclusive deep inelastic scattering data from HERA and consider two different starting points $x_{0}$ for the evolution. We find better agreement with the experimental data for a smaller $x_{0}$. This work paves the way for future studies involving global fits of the dipole amplitude at leading order and beyond.
@@ 1 Introduction
Mapping the partonic structure of protons and nuclei in the high-energy regime is a foremost challenge that provides the primary scientific motivation for the future Electron-Ion Colliders Accardi and others (2016); Abdul Khalek and others (2022); Anderle and others (2021). At a very small Bjorken-$x$ variable, the density of gluons within a hadron grows rapidly, leading to the gluon saturation phenomenon where non-linear QCD dynamics become dominant. The Color Glass Condensate (CGC) effective field theory Iancu and Venugopalan (2003); Gelis et al. (2010) provides a powerful framework for describing this dense gluonic matter. Central to the CGC is the dipole–target scattering amplitude, which encodes the essential information about the nonperturbative scattering off the gluonic target. The energy evolution of the dipole amplitude is governed by perturbative evolution equations, most notably the Balitsky–Kovchegov (BK) equation Balitsky (1996); Kovchegov (1999). However, the starting point for this evolution—the initial condition at a moderately small $x_{0}$—is fundamentally nonperturbative and must be determined from experimental data.
@@
A convenient initial condition for the high-energy evolution is given by the McLerran–Venugopalan (MV) model McLerran and Venugopalan (1994b, a, c), which allows for a description of the dipole amplitude in terms of a few free parameter that can be extracted from the experimental data. The MV model, and its generalizations, have then be used in many successful extractions at both leading order Albacete et al. (2009); Lappi and Mäntysaari (2013); Albacete et al. (2011); Ducloué et al. (2020); Beuf et al. (2020); Dumitru et al. (2023); Casuga et al. (2024); Hänninen et al. (2025) and next-to-leading order Hänninen et al. (2023); Casuga et al. (2025) using the inclusive deep inelastic scattering (DIS) data from HERA Aaron and others (2010); Abramowicz and others (2015). The extracted parameters contain the fundamental information about the target structure, including the saturation scale $Q_{s}$ that describes the onset of the gluon saturation phenomenon.
@@
However, to determine the initial condition from the experimental data, one has to run the BK evolution for a wide range of parameter values. Due to the nonlinear nature of the BK evolution, its evaluation is much more demanding than the linear BFKL Lipatov (1976); Fadin et al. (1975); Kuraev et al. (1976, 1977); Balitsky and Lipatov (1978) and DGLAP Dokshitzer (1977); Gribov and Lipatov (1972a, b); Altarelli and Parisi (1977) evolution equations, making the BK evolution the bottleneck for numerical studies even at leading order. At higher orders in perturbation theory, the cross section itself becomes also numerically demanding due to the large amount of integrals over transverse coordinates that are typical for the dipole picture used in small-$x$ calculations. For these reasons, it becomes important to study efficient methods for evaluating these quantities with different models for the dipole amplitude.
To overcome this limitation, we turn to the rapidly advancing field of machine learning and neural-network emulation. We employ the transformer architecture, originally introduced for sequence modelling with self-attention Vaswani et al. (2017). Transformers are widely used to capture long-range dependencies in sequential data, including long-horizon time-series forecasting Zhou et al. (2021); Lim et al. (2021). In high-energy physics, transformer architectures have also been successfully applied to learn correlations in variable-length particle sets, e.g., for jet tagging Qu et al. (2022) and symmetry-aware event reconstruction tasks such as jet–parton assignment Shmakov et al. (2022); Fenton et al. (2022).
Motivated by these developments, we treat the model parameters together with the kinematic variables as an ordered sequence of inputs. In this way, the transformer’s self-attention mechanism Vaswani et al. (2017) can learn intricate, non-local correlations in the BK-evolved dipole amplitude and the corresponding DIS dataset (such as those from HERA) without imposing a rigid functional ansatz.
From a functional viewpoint, the BK evolution is defined as a non-linear differential equation that resums logarithms of energy. For a given set of initial-condition parameters and a given kinematic point $(r,x)$, it returns the dipole amplitude $N(r,x)$ that has been evolved starting from an initial condition at $x_{0}>x$. In practical fits one must evaluate this map for many parameter choices and across a wide kinematic domain spanning many orders of magnitude in both $r$ and $x$. As a result, the relevant target function is strongly multi-scale and exhibits correlated variations between the initial-condition parameters and the kinematic variables.
The self-attention mechanism in transformers provides a flexible way to model such cross-dependencies: by allowing each input component (parameters and kinematics) to interact with all others, the network can learn global correlations that are difficult to capture with rigid low-dimensional ansätze. This perspective is closely related to recent developments in scientific machine learning, where attention-based architectures have been explored for surrogate/operator learning of complex non-linear systems Hao et al. (2023); Li et al. (2023).
To our knowledge, this paper presents the first determination of the small-$x$ dipole amplitude using this transformer-based emulator strategy. Moreover, we consider different starting points for the BK evolution, enabling us to study potential biases in the fitted initial condition and to assess the applicability of BK evolution when starting from more moderate values of Bjorken $x$. It should be emphasized that the present analysis is restricted to leading-order accuracy, and our primary aim is to demonstrate the methodology rather than to obtain the most precise phenomenological fit.
An alternative machine-learning approach based on Gaussian-process emulators has been successfully applied in small-$x$ studies before Casuga et al. (2024, 2025), providing flexible surrogate models with quantified uncertainties. In this work, we employ a transformer-based approach, which can naturally accommodate larger training sets, capture correlations across the parameter space, and demonstrates stable performance for learning such highly non-linear multi-variable mappings Hao et al. (2023); Li et al. (2023). This makes it well suited for global analyses requiring high precision over a wide kinematic range.
@@ 2.1 Inclusive DIS in the dipole picture
At small $Q^{2}$ region, where contributions from $Z$ boson exchange are neglected, inclusive DIS is described by the structure functions $F_{2}(x,Q^{2})$ and $F_{L}(x,Q^{2})$, which are related to the experimental reduced cross section $\sigma_{r}$ by
@@
In the small-$x$ regime, these cross sections can be calculated within the dipole picture Kovchegov and Levin (2013), where the interaction factorizes into the splitting of the virtual photon into a quark–antiquark pair ($q\bar{q}$), followed by the scattering of this color dipole off the proton target. The total cross section is given by an integral over the dipole’s transverse size $\mathbf{r}$, its impact parameter $\mathbf{b}$, and the quark’s longitudinal momentum fraction $z$, which reads
@@
and the argument $\overline{Q}$ of the modified Bessel functions $K_{0}$ and $K_{1}$ is $\overline{Q}^{2}\equiv z(1-z)Q^{2}$. Here, we have ignored quark mass corrections. In many dipole-model phenomenological studies, quark-mass threshold effects are approximately accounted for by a rescaling of the Bjorken variable, $\tilde{x}=x\,(1+4m_{f}^{2}/Q^{2})$ (the original GBW prescription) Golec-Biernat and Wusthoff (1998). In this work, since we adopt massless photon wave functions at leading order, we do not apply this $x$-rescaling.
@@
and introducing a single normalization parameter, $\sigma_{0}/2$, which has the interpretation as the proton transverse area  Lappi and Mäntysaari (2013); Albacete et al. (2011, 2009). This parameter effectively sets the total strength of the interaction by replacing the impact parameter integral in the cross-section calculation. It is a free parameter in our model, constrained by the fit to HERA data. The integrated dipole amplitude $N(r,x)$ now only depends on the dipole size $|r=\absolutevalue{{\mathbf{r}}}$ and the Bjorken $x$ variable. Due to rotational symmetry, there is no dependence on the angle of the vector $\mathbf{r}$} after integrating over the impact parameter.
@@ 2.2 Balitsky–Kovchegov evolution
The energy dependence of the dipole amplitude, $N(r,x)$, is governed by the BK evolution equation. This equation resums quantum corrections proportional to large logarithms of energy, typically written in terms of the rapidity variable $Y=\ln(1/x)$, and incorporates the non-linear effects that lead to gluon saturation. For a dipole with quark and antiquark at transverse positions ${\mathbf{r}}_{0}$ and ${\mathbf{r}}_{1}$, the equation reads
@@
The interaction kernel, $K_{\text{BK}}$, describes the splitting of the parent dipole. For phenomenological accuracy, it is crucial to include running coupling corrections. Following the Balitsky prescription Balitsky (2007), the kernel is given by
@@
Here the position-space strong coupling constant, $\alpha_{s}(r^{2})$, is taken at leading order with $n_{f}=3$ active quark flavors as Lappi and Mäntysaari (2013)
@@
Here, $\Lambda_{\text{QCD}}=$0.241\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$\$ and $C^{2}$ is a parameter that sets the scale of the argument of the logarithm, which must be determined from fits to data Lappi and Mäntysaari (2013); Casuga et al. (2024). For large values of the dipole size $r$, we freeze $\alpha_{s}$ to the value $\alpha_{s,\text{max}}=0.7$.
@@ 3.1 Sampling and solving the BK equation
The BK equation (8) describes the evolution of the dipole amplitude $N(r,x)$ with respect to the collision energy. The evolution is uniquely specified by an initial condition at $x_{0}$, and to assess the dependence on the initial rapidity scale we consider two different choices $x_{0}=0.01$ and $x_{0}=0.05$. For the initial condition, we use a generalization of the MV model considered in Ref. Lappi and Mäntysaari (2013)
@@
which is a function of the initial saturation scale $Q_{s0}$, the anomalous dimension $\gamma$, and an infrared regulator $e_{c}$. To efficiently map this parameter space, we use Latin hypercube sampling (LHS) to generate 10 00010\,000 parameter sets for $(Q_{s0},\gamma,e_{c},C^{2})$. For each set, the running-coupling BK equation with the Balitsky kernel is solved using a numerical Julia implementation developed for this work. The solution is tabulated on a two-dimensional grid spanning transverse sizes $10^{-6}\leq r/\mathrm{GeV}^{-1}\leq 10^{2}$ and rapidities $0\leq\ln(x_{0}/x)\leq 16$, which covers the relevant HERA kinematics Abramowicz and others (2015). This procedure yields a dataset of approximately $4.7\times 10^{7}$ values of $N(r,x)$, which serves as the training data for the neural network model.
For the neural-network inputs and diagnostics, we work in log-transformed variables for the two scale parameters, using $\ln Q_{s0}$ and $\ln C^{2}$, which improves numerical conditioning across the wide dynamic ranges. The size of the training library (10 000 configurations) is chosen as a practical balance between the dimensionality of the parameter space, the target emulator accuracy, and the computational cost of generating BK solutions. In general, using substantially fewer configurations reduces the coverage of the parameter space and tends to increase the emulator errors, while increasing the library size improves accuracy at the expense of additional BK-evolution time.
The BK solutions are tabulated as $N(r,x)$ on a two-dimensional output grid summarized in Table 1. The $r$ grid is logarithmic (uniform in $\ln r$), while the $x$ output points are uniformly spaced in $\ln(x_{0}/x)$ with step $\Delta\ln(x_{0}/x)=0.2$. Here $\Delta\ln(x_{0}/x)$ denotes only the output spacing. To reduce the size of the training dataset, we keep only samples with $\ln N(r,x)\geq-10$ when forming the training set; this cut is applied only in data preprocessing and does not affect the BK evolution itself.
Quantity Value
$r$ grid logarithmic (uniform in $\ln r$)
$r_{\min}$ $\mathrm{GeV}^{-1}$] $10^{-6}$
$r_{\max}$ $\mathrm{GeV}^{-1}$] $10^{2}$
$N_{r}$ $150$
$x$ output grid uniform in $\ln(x_{0}/x)$
$\ln(x_{0}/x)_{\min}$ 0
$\ln(x_{0}/x)_{\max}$ $16$
$\Delta\ln(x_{0}/x)$ $0.2$
Table 1: Output grid used to store the BK solutions. The $x$ points are logarithmically spaced, corresponding to uniform spacing in $\ln(x_{0}/x)\in[0,16]$ with $\Delta\ln(x_{0}/x)=0.2$.
@@ 3.2 Neural network emulation of the BK amplitude
Figure 1: Left: Two dimensional histogram comparing the exact and transformer emulated values of $\ln N(r,x)$ for the BK dipole amplitude. The color scale indicates the density of samples, and the dashed line represents exact agreement. The mean and median relative errors on the validation set are $0.09\%$ and $0.05\%$, respectively. Right: Distribution of the signed relative errors, zoomed to $\pm 1\%$. This highlights that the vast majority of predictions fall within a few per mille of the true values, demonstrating the emulator’s high accuracy. We find that $99.978\%$ of the validation points fall within the $\pm 1\%$ interval.
Directly solving the BK equation repeatedly across the parameter space would be computationally expensive. To circumvent this we train a neural network to emulate the functional dependence of $\ln N(r,x)$ on the variables $(r,x)$ and the four parameters of the BK setup. Owing to its ability to capture correlations in an ordered sequence of input variables, we adopt a transformer architecture. The network takes as input the 6-dimensional vector $(\ln Q_{s0},\,\allowbreak e_{c},\,\allowbreak\gamma,\,\allowbreak\ln C^{2},\,\allowbreak\ln(r/\mathrm{GeV}^{-1}),\,\allowbreak\ln(x_{0}/x))$. A single input vector corresponds to one training sample, i.e. one tabulated BK value at a specific grid point $(r,x)$ for a specific parameter set $\bm{\theta}=(Q_{s0},e_{c},\gamma,C^{2})$. Concretely, the first four entries specify the BK setup (the initial-condition parameters $\bm{\theta}$), and the last two entries specify the kinematics at which the amplitude is evaluated. The full training dataset is constructed by looping over all sampled parameter sets and, for each of them, over all $(r,x)$ points on the numerical output grid used to tabulate the BK solutions. Each scalar is mapped by a shared linear projection to form a sequence of six tokens of dimension $d_{\rm model}=128$ (see App. A for terminology). A sinusoidal positional encoding is added to the sequence, which is then processed by a transformer encoder with four layers, eight attention heads, feed-forward width 512, dropout 0.1, and pre–normalization. The token representations are mean-pooled and passed to a two-layer multilayer perceptron (MLP) head Gardner and Dorling (1998) that outputs a raw logit function Wei et al. (2022). Applying a sigmoid to this logit yields the dipole amplitude $N(r,x)$, automatically enforcing the unitarity bound $0\leq N\leq 1$. For diagnostics we also use $\log N$ after the sigmoid.
@@
Inputs are standardized using the StandardScaler implementation from the Scikit-learn library Pedregosa et al. (2011). The model is trained using the AdamW optimizer Loshchilov and Hutter (2017) with learning rate $3\times 10^{-4}$, weight decay $10^{-4}$, cosine annealing schedule to $3\times 10^{-6}$, batch size $4096$, and early stopping (patience $20$). Automatic mixed-precision training Micikevicius et al. (2017) and gradient clipping (max-norm $1.0$) are applied. $90\%$ of the generated data is used for training and the remainder for validation.
Figure 2: The emulated dipole amplitude $N(r,x)$ (lines) is compared against the exact BK solution (points) for a representative out-of-training-sample parameter set at three different rapidities, $x=10^{-3}$, $10^{-5}$, and $10^{-7}$. The excellent agreement in both log-linear (left) and log-log (right) scales across the entire range of $r$ demonstrates the model’s strong generalization performance on a held-out parameter set. The chosen parameters are $Q_{s0}^{2}=0.07$ GeV2, $\gamma=1.01$, $e_{c}=24.68$, $C^{2}=4.65$, and $x_{0}=0.01$, corresponding roughly to physical values found in previous fits Lappi and Mäntysaari (2013); Casuga et al. (2024).
The trained emulator exhibits excellent performance, achieving mean and median relative errors of $0.09\%$ and $0.05\%$, respectively, on the validation set. This high accuracy is visually demonstrated in Fig. 1. The left panel shows a strong one-to-one correlation between the emulated and exact values of $\ln N(r,x)$ across many orders of magnitude. The distribution of relative errors, shown in the right panel, is sharply peaked at zero, confirming that the vast majority of predictions are accurate to within a few per mille. The slight asymmetry in the 2D histogram reflects the non-uniform distribution of validation samples across the multi-scale $(r,x)$ domain and parameter space, and is amplified by binning in the density plot. The relative errors quantify the emulator accuracy with respect to the tabulated BK targets on a held-out validation set, and should not be interpreted as the intrinsic numerical uncertainty of the BK solver.